OpenStudy (anonymous):

If a wheel with a radius of 80 inches spins at a rate of 50 revolutions per minute, find the approximate linear velocity in miles per hour/ Use 3.14 for pi. Thank you!

OpenStudy (welshfella):

first work out the circumference of the wheel can you do that?

OpenStudy (welshfella):

Circumference = 2 * 3.14 * radius

OpenStudy (anonymous):

yes, 2pi r= 502.4

OpenStudy (welshfella):

= 2* 3.14 * 80 = ?

OpenStudy (welshfella):

50 revolutions per minute = 50 * 502.4 ins per minute

OpenStudy (anonymous):

that's how I get stuck! it's inches to miles and minutes to hours... I've tried it about 5 different ways before I posted. (It's multiple choice, so I know my answers are wrong when they aren't there!)

OpenStudy (welshfella):

I'm not at my best today - i just said divide by 60 instead of multiply its 50*502.4 * 60 inches per hour then as there are 5280 * 12 inches ina mile mph = 50*502.4 * 60 ------------- 5280*12

OpenStudy (welshfella):

= 23.79 mph ?

OpenStudy (welshfella):

is that a choice?

OpenStudy (anonymous):

ahhhh! Thanks found my error. I was forgetting the 50 back in the original formula. Early for me too, apparently.

OpenStudy (welshfella):

lol ok